# Coherent High-Harmonic Generation with Laser-Plasma Beams

### v2

@@ Abstract
Active energy compression allows the LPA beam reaching extremely low levels of energy spread and jitter, down to $\sim 10^{-5}$, employing conventional radiofrequency (RF) accelerating cavities for active energy compression [3]. This scheme is envisioned for the PETRA IV plasma injector [1] to enable clean and efficient injection in the storage ring. Once filled, the storage ring needs to be topped up only once every few minutes to restore a small fraction ($\sim 1$) of the total charge lost due to Touschek and residual gas scattering  [^1] Contemporary and planned 4th generation light sources, such as PETRA IV feature large $\gtrsim 10$ h beam lifetimes.. This opens a window of opportunity for additional applications of an LPA injector that could take advantage of high quality electron beams, when they are not required for the storage ring. Thanks to their small energy spread after energy compression the LPA beams require a relatively small energy modulation to access high $\sim 100$-th harmonics of the seed wavelength, while a high power laser for the seeding is naturally present in the LPA setup. Thus, the LPA electron beams can produce stable, narrow-band coherent radiation at high harmonics of the driving laser pulse, achieving wavelengths as small as 10 nm for an 800 nm seed.
Figure 1(a) shows the proposed setup, based on a 500 MeV LPA injector described in Ref. [3]. It consists of an LPA followed by a quadrupole triplet to capture the electron beam from the plasma cell, a chicane with sextupoles for chromaticity correction, a large stretcher chicane, and finally an RF cavity for suppressing energy spread and correcting energy deviations. The injector beamline drastically reduces the relative energy spread of the beam from $\sim 10^{-2}$ to $<10^{-4}$ [Fig. 1(b,c)].
@@
The energy-compressed beam then enters the first undulator where it co-propagates with a seed laser pulse that is generated by splitting a small fraction of the drive pulse. The laser field imprints a modulation of the beam energy with a spatial period of the laser wavelength [Fig. 1(d)]. After passing through a dispersive section the energy modulation translates into microbunching with the periodicity of the modulation wavelength [Fig. 1(e)]. The resulting microbunches have a peak current far exceeding that of the initial bunch. And finally, the resulting train of short microbunches radiates in the second undulator at the high $n\gg 1$ harmonic of the seed laser wavelength $\lambda_{l}$.
@@
where $e$ is the elementary charge, $\gamma=E_{0}/mc^{2}$ is the mean Lorentz factor of the beam, $P_{l}$ is the modulator laser power, $w_{0}$ the laser beam waist size, $Z_{0}\approx 377~{}\Omega$ the vacuum impedance, $L_{u}=\lambda_{u}N_{w}$ the undulator length (with $\lambda_{u}$ the undulator period and $N_{w}$ the number of periods), $K$ the undulator parameter, and $A_{JJ}=J_{0}(\chi)-J_{1}(\chi)$, $J_{n}$ is the $n$-th Bessel function of the first kind, and $\chi=K^{2}/(4+2K^{2})$. The undulator $K$ has to be chosen such that its resonance wavelength is tuned to the seed laser:
@@
A few additional effects might affect the idealized picture presented above. First, as the bunch passes the undulator, the undulator generates an additional energy-position correlation: $R_{56,u}=2\lambda N_{w}.$ The additional $R_{56}$ created by the modulator affects the optimum condition in Eq. (5), reducing the $R_{56}$ that needs to be created by the chicane. In the radiator it affects the microbunching by displacing the beam from the optimum condition set by Eq. (5) and thus limiting the distance over which the high bunching factor remains. We can estimate this distance by applying the condition $R_{56,u}/R_{56}\lesssim\Delta_{a}$, where $\Delta_{a}$ is the relative full-width half-maximum of the bunching factor peak. For the 80-th harmonic it is about 0.1, while for the 27-th harmonic it is about 0.2 (Fig. 2).
@@
As an example, we consider the 500 MeV energy beam produced by the injector from Ref. [3], shown in Fig. 1(b,c). Table 1 summarizes the most relevant beam parameters after the energy compression. For coherent high-harmonic generation, we consider two scenarios. In case I, the beam is modulated by a laser pulse with the wavelength of the LPA driver: $\lambda_{l}=800$ nm, obtained by picking up a fraction of the drive laser. In case II, the modulation is performed by a laser pulse with three times shorter wavelength: $\lambda_{l}=270$ nm, which can be obtained by frequency-conversion of a fraction of the driver laser [6]. In both cases, we target a final radiation wavelength of 10 nm, which corresponds to the 80-th harmonic of 800 nm and the 27-th of 270 nm. According to Fig. 2, in order to achieve a significant bunching factor at those harmonics, we select an energy modulation of $A=100$ and $A=50$, which corresponds to an absolute energy modulation amplitude of 2.50 MeV and 1.25 MeV, for cases I and II, respectively. This modulation requires a laser peak power of 2.3 (1.8) GW and 48 (36) mJ energy for case I (II), according to Eq. (1). This is a small fraction of the 2.45 J pulse used to produce and accelerate the electron beam in the LPA.
Table 1: Electron beam parameters
@@
We used Ocelot [19] to track the electron beam through the harmonic generation beamline [Fig. 1 (a)] and simulate its density modulation. Beamline elements were modelled by linear and second-order transfer maps. Figure 4 shows the betatron beam sizes throughout the whole beamline. To minimize the beam divergence, these were kept approximately constant at $\lesssim 100~{}\upmu$m through the modulator plus chicane in both horizontal and vertical planes. The energy modulation process in the modulator was simulated accounting for the finite size and diffraction of the laser pulse. In order to mitigate these effects a sufficiently large value for the laser waist, $w_{0}=1$ mm was selected. Table 2 summarizes the undulator parameters. Note that the magnetic field of the undulator was decreased in case II to account for the smaller $\lambda_{l}$.
@@
Due to the high number of particles required to properly resolve the coherent radiation, we limited the simulation to the central slice of the beam where we track every single electron. Thus, a total of $4\times 10^{6}$ particles were used to generate a flat-top beam slice of length 16 $\upmu$m and current 12 A at the entrance of the modulator. Figure 1(d) presents the longitudinal phase space of a central beam slice right after the modulator. The following magnetic chicane was tuned to maximize the bunching factor of the beam at the middle of the radiator [Fig. 1(e)]. For case I (II), this resulted in a chicane with $R_{56}=17~{}(13)~{}\upmu$m, just a 70% of the value calculated from Eq. (5). The additional compression was given passively by the combined effect of the modulator and radiator, and the free drifts in between. In case I (II), the central slice spans a total of 20 (60) microbunches. The Courant-Snyder parameters of the beam slice differ slightly from those of the whole beam (Table 1333In particular, the normalized emittances are 2.1 and 0.4 $\mathrm{\upmu m}$ in the horizontal and transverse planes, respectively, slightly smaller than the projected ones [see Fig. 1(c)]. This resulted in smaller transverse sizes (rms) of the slice: $\sigma_{x}=73~{}\upmu$m and $\sigma_{y}=44~{}\upmu$m at the center of the radiator.
@@
The radiation field emitted by the beam in the final undulator was calculated using the code Synchrad [26]. This code computes the coherent sum of the radiation field produced individually by the particles composing the beam, using the Fourier transformed Lienard-Wiechert potentials [27]. The particle trajectories through the radiator were calculated analytically with a resolution of 64 points per undulator period, assuming a perfectly sinusoidal magnetic field and neglecting collective effects and radiation energy losses. The particle trajectories were fed into Synchrad to calculate the radiation field. Figure 5 shows the peak brilliance of the radiation for the two considered cases. In logarithmic scale, we can see how the coherent radiation is about five orders of magnitude greater than the incoherent signal, and characterized by the presence of narrow peaks at the harmonics of the seed. As expected, the maximum peak brilliance is reached where the fundamental radiation wavelength of the undulator is resonant with one of the harmonics, which for case I (II) is 1.7 (4.5) ×$\times 10^{25}~{}\mathrm{photons/s/mm^{2}/mrad^{2}/0.1\%bw}$. Figure 5 also shows an inset in linear scale where we observe that the spectrum is strongly dominated by the resonant peak, featuring a relative rms bandwidth of $1.8\times 10^{-4}$ and $3.9\times 10^{-4}$ for cases I and II, respectively. The average radiated power for case I (II) yields 1.5 (6.3) kW. To estimate the total radiated energy, we simply integrate the result of the central slice through the longitudinal profile of the whole beam. Assuming a Gaussian temporal profile with $2.7$ ps rms (Table 1), we obtain 0.01 (0.04) $\upmu$J for case I (II).
@@
In conclusion, with the help of the energy compression technique modern LPAs are approaching a point where they will be capable of delivering electron beams with low energy spread and jitter. We are proposing a scheme that takes advantage of this low energy spread to generate very high, $\sim 100$-th harmonics of the driver laser. Our estimates have shown that using a small, $\sim 1\%$ fraction of the driver laser pulse to modulate the electron beam energy one can form a train of few thousands microbunches within the full width at half maximum of the initial electron bunch. This is possible thanks to the low, $10^{-5}$-level energy spread of the electron beam. The resulting short microbunches then allow accessing high harmonics of the modulator wavelength in the radiator undulator, producing $\sim 1$-ps-long bursts of radiation. The radiation wavelength can be easily tuned by selecting the desired harmonic of the seed $\lambda_{l}$, adjusting the radiator $K$ parameter or the beam energy. The distance between adjacent harmonics is about $\lambda_{l}/n^{2}\approx 0.1$ nm for the 80-th harmonic of 800 nm. Similar radiation power can be generated at each wavelength, because the bunching factor depends weakly on $n$ for high $n$’s (Fig. 3). Since the width of the resulting radiation is determined by the spectrum of the beam current, the long microbunch trains would generate extremely narrow spectral lines, as seen in our numerical simulation. Ultimately, one can aim at achieving the width as small as $\sim 10^{-5}$. In practice, the spectral width and phase noise of the seed laser pulse and the nonlinear energy chirp, remaining in the electron beam after the energy compression, will broaden the spectrum. Together these effects might worsen the relative width to $\sim 10^{-4}$ level. The narrow bandwidth might find applications in spectroscopy where a high spectral resolving power is required. The proposed scheme has a potential for scalability to higher beam energies, which can help mitigating detrimental effects of space charge and beam divergence. Since the radiation wavelength is defined by the harmonics of the seed wavelength, a higher beam energy may not always result in a higher photon energy. Still, it may enable a wider range of undulator parameters such as a longer period and higher strength, allowing for additional tunability of the resulting radiation.

### v3

@@ I Introduction
Active energy compression allows the LPA beam reaching extremely low levels of energy spread and jitter, down to $\sim 10^{-5}$, employing conventional radiofrequency (RF) accelerating cavities for active energy compression [3]. This scheme is envisioned for the PETRA IV plasma injector [1] to enable clean and efficient injection in the storage ring. Once filled, the storage ring needs to be topped up only once every few minutes to restore a small fraction ($\sim 1$) of the total charge lost due to Touschek and residual gas scattering  [^1] Contemporary and planned 4th generation light sources, such as PETRA IV feature large $\gtrsim 10$ h beam lifetimes.. This opens a window of opportunity for additional applications of an LPA injector that could take advantage of high quality electron beams, when they are not required for the storage ring. Thanks to their small energy spread after energy compression the LPA beams require a relatively small energy modulation to access high $\sim 100$-th harmonics of the seed wavelength, while a high power laser for the seeding is naturally present in the LPA setup. Thus, the LPA electron beams can produce stable, narrow-band radiation at high harmonics of the driving laser pulse, achieving wavelengths as small as 10 nm for an 800 nm seed.
Figure 1(a) shows the proposed setup, based on a 500 MeV LPA injector described in Ref. [3]. It consists of an LPA followed by a quadrupole triplet to capture the electron beam from the plasma cell, a chicane with sextupoles for chromaticity correction, a large stretcher chicane, and finally an RF cavity for suppressing energy spread and correcting energy deviations.
@@
The injector beamline drastically reduces the relative energy spread of the beam from $\sim 10^{-2}$ to $<10^{-4}$ [Fig. 1(b,c)]. The energy-compressed beam then enters the first undulator where it co-propagates with a seed laser pulse that is generated by splitting a small fraction of the drive pulse. The laser field imprints a modulation of the beam energy with a spatial period of the laser wavelength [Fig. 1(d)]. After passing through a dispersive section the energy modulation translates into microbunching with the periodicity of the modulation wavelength [Fig. 1(e)]. The resulting microbunches have a peak current far exceeding that of the initial bunch. And finally, the resulting train of short microbunches radiates in the second undulator at the high $n\gg 1$ harmonic of the seed laser wavelength $\lambda_{l}$.
II Analytical estimates
@@
where $e$ is the elementary charge, $\gamma=E_{0}/mc^{2}$ is the mean Lorentz factor of the beam, $P_{l}$ is the modulator laser power, $w_{0}$ the laser beam waist size, $Z_{0}\approx 377~{}\Omega$ the vacuum impedance, $L_{u}=\lambda_{u}N_{w}$ the undulator length (with $\lambda_{u}$ the undulator period and $N_{w}$ the number of periods), $K$ the undulator parameter, and $A_{JJ}=J_{0}(\chi)-J_{1}(\chi)$, $J_{n}$ is the $n$-th Bessel function of the first kind, and $\chi=K^{2}/(4+2K^{2})$. The undulator $K$ has to be chosen such that its resonance wavelength is tuned to the laser:
@@
III Additional considerations
A few additional effects might affect the idealized picture presented above. First, as the bunch passes the undulator, the undulator generates an additional energy-position correlation: $R_{56,u}=2\lambda N_{w}$, where $\lambda$ stands for the photon wavelength, $\lambda_{l}$ in the modulator and $\lambda_{l}/n$ in the radiator. The additional $R_{56}$ created by the modulator affects the optimum condition in Eq. (5), reducing the $R_{56}$ that needs to be created by the chicane. In the radiator it affects the microbunching by displacing the beam from the optimum condition set by Eq. (5) and thus limiting the distance over which the high bunching factor remains. We can estimate this distance by applying the condition $R_{56,u}/R_{56}\lesssim\Delta_{a}$, where $\Delta_{a}\approx 1.8/n^{2/3}$ is the relative full-width half-maximum of the bunching factor peak. For the 80-th harmonic it is about 0.1, while for the 27-th harmonic it is about 0.2 (Fig. 2).
@@ III Additional considerations
The bunching factor can also be affected by the coherent synchrotron radiation (CSR). CSR effects in the beam delivery line before the energy compression have been studied previously [3] and found to have no significant effect on the beam quality thanks to a rapid decompression of the beam. Let us estimate the magnitude of the CSR kick after the energy modulation using a simple steady-state 1D model [25]: $W_{\mathrm{CSR}}\approx N_{b}m_{e}c^{2}r_{e}\kappa^{2/3}\sigma_{z,\mathrm{mb}}^{-4/3}$, where $\kappa=0.042~{}$m-1 is the bending curvature and $\sigma_{z,\mathrm{mb}}$ is the rms length of a microbunch. In the beginning of the chicane $\sigma_{z,\mathrm{mb}}\sim\lambda_{l}$ and $I_{\mathrm{mb}}\sim 5$ A, resulting in negligible $W_{\mathrm{CSR}}\sim 5$ keV/m. Approaching the end of the chicane $\sigma_{z,\mathrm{mb}}$ reaches $\sim\lambda_{l}/n$ and $I_{\mathrm{mb}}\sim 400$ A, the CSR becomes significant: $W_{\mathrm{CSR}}\approx 1.7~{}$MeV/m. But since the dipole is only 15 cm long, this results in only $5\times 10^{-4}$ relative energy variation – a small figure compared to the laser modulation amplitude.
@@
IV Numerical examples
As an example, we consider the 500 MeV energy beam produced by the injector from Ref. [3], shown in Fig. 1 (a). First, the beam is obtained through Bayesian optimization of the LPA stage using the particle-in-cell code FBPIC [18] and the optimization software Optimas [26, 27] [Fig. 1(b)]. Then, the beam is tracked through the energy compression and harmonic generation parts of the beamline in the Ocelot [19] code, which uses up to second-order transfer maps to simulate the beamline elements. CSR effects through the dispersive sections of the beamline (the chicanes) are computed by Ocelot using a projected 1D model [28]. Figure 1(c) shows the beam after the energy compression and Tab. 1 summarizes its most relevant parameters at this point. Figures 1(d) and 1(e) show the result of the tracking through the energy modulator and the subsequent dispersive section for a central slice of the beam. Finally, the coherent radiation emitted by the beam in the radiator is computed using the code Synchrad [29]. Throughout this start-to-end simulation chain, the same macroparticles representing the simulated beam are transported from code to code transforming to the corresponding coordinate system.
Table 1: Beam parameters after energy compression.
@@ IV Numerical examples
For coherent high-harmonic generation, we consider two scenarios. In case I, the beam is modulated by a laser pulse with the wavelength of the LPA driver: $\lambda_{l}=800$ nm, obtained by picking up a fraction of the drive laser. In case II, the modulation is performed by a laser pulse with three times shorter wavelength: $\lambda_{l}=270$ nm, which can be obtained by frequency-conversion of a fraction of the driver laser [6]. In both cases, we target a final radiation wavelength of 10 nm, which corresponds to the 80-th harmonic of 800 nm and the 27-th of 270 nm. According to Fig. 2, in order to achieve a significant bunching factor at those harmonics, we select an energy modulation of $A=100$ and $A=50$, which corresponds to an absolute energy modulation amplitude of 2.50 MeV and 1.25 MeV, for cases I and II, respectively. This modulation requires a laser peak power of 2.3 (1.8) GW and 48 (36) mJ energy for case I (II), according to Eq. (1). This is a small fraction of the 2.45 J pulse used to produce and accelerate the electron beam in the LPA. Table 2 summarizes the undulator parameters. Note that the magnetic field of the undulator was decreased in case II to account for the smaller $\lambda_{l}$.
@@
Ocelot is also used to simulate the energy modulation process in the modulator accounting for the finite size and diffraction of the laser pulse. In order to mitigate these effects a sufficiently large value for the laser waist, $w_{0}=1$ mm is selected. Figure 4 shows the betatron beam sizes throughout the whole beamline. To minimize the beam divergence, these are kept approximately constant at $\lesssim 100~{}\upmu$m through the modulator plus chicane in both horizontal and vertical planes.
Due to the high number of particles required to properly resolve the coherent radiation, we limit the simulation to the central slice of the beam where we track every single electron. Thus, a total of $4\times 10^{6}$ particles are used to generate a flat-top beam slice of length 16 $\upmu$m and current 12 A at the entrance of the modulator. Figure 1(d) presents the longitudinal phase space of a central beam slice right after the modulator. The following magnetic chicane is tuned to maximize the bunching factor of the beam at the middle of the radiator [Fig. 1(e)]. For case I (II), this results in a chicane with $R_{56}=17~{}(13)~{}\upmu$m, just a 70% of the value calculated from Eq. (5). The additional compression is given passively by the combined effect of the modulator and radiator, and the free drifts in between. In case I (II), the central slice spans a total of 20 (60) microbunches. The Courant-Snyder parameters of the beam slice differ slightly from those of the whole beam (Table 1333In particular, the normalized emittances are 2.1 and 0.4 $\mathrm{\upmu m}$ in the horizontal and transverse planes, respectively, slightly smaller than the projected ones [see Fig. 1(c)]. This results in smaller transverse sizes (rms) of the slice: $\sigma_{x}=73~{}\upmu$m and $\sigma_{y}=44~{}\upmu$m at the center of the radiator.
@@
The radiation field emitted by the beam in the final undulator is calculated using the code Synchrad [29]. This code computes the coherent sum of the radiation field produced individually by the particles composing the beam, using the Fourier transformed Lienard-Wiechert potentials [31]. To compare with the case of a non-modulated beam, we also use Synchrad to compute the incoherent spectrum as the sum of the radiation intensities of individual electrons. The particle trajectories through the radiator are calculated analytically with a resolution of 64 points per undulator period, assuming a perfectly sinusoidal magnetic field and neglecting collective effects and radiation energy losses. The particle trajectories are fed into Synchrad to calculate the radiation field. Figure 5 shows the peak brilliance of the radiation for the two considered cases. In logarithmic scale, we can see how the coherent radiation is about five orders of magnitude greater than the incoherent signal, and characterized by the presence of narrow peaks at the harmonics of the seed. As expected, the maximum peak brilliance is reached where the fundamental radiation wavelength of the undulator is resonant with one of the harmonics, which for case I (II) is 1.7 (4.5) ×$\times 10^{25}~{}\mathrm{photons/s/mm^{2}/mrad^{2}/0.1\%bw}$. Figure 5 also shows an inset in linear scale where we observe that the spectrum is strongly dominated by the resonant peak, featuring a relative rms bandwidth of $1.8\times 10^{-4}$ and $3.9\times 10^{-4}$ for cases I and II, respectively. The average radiated power for case I (II) yields 1.5 (6.3) kW. To estimate the total radiated energy, we simply integrate the result of the central slice through the longitudinal profile of the whole beam. Assuming a Gaussian temporal profile with $2.7$ ps rms (Table 1), we obtain 0.01 (0.04) $\upmu$J for case I (II).
@@
V Conclusion
With the help of the energy compression technique modern LPAs are approaching a point where they will be capable of delivering electron beams with low energy spread and jitter. We are proposing a scheme that takes advantage of this low energy spread to generate very high, $\sim 100$-th harmonics of the driver laser. Our estimates have shown that using a small, $\sim 1\%$ fraction of the driver laser pulse to modulate the electron beam energy one can form a train of few thousands microbunches within the full width at half maximum of the initial electron bunch. This is possible thanks to the low, $10^{-5}$-level energy spread of the electron beam. The resulting short microbunches then allow accessing high harmonics of the modulator wavelength in the radiator undulator, producing $\sim 1$-ps-long bursts of radiation. The radiation wavelength can be easily tuned by selecting the desired harmonic of the seed $\lambda_{l}$, adjusting the radiator $K$ parameter or the beam energy. The distance between adjacent harmonics is about $\lambda_{l}/n^{2}\approx 0.1$ nm for the 80-th harmonic of 800 nm. Similar radiation power can be generated at each wavelength, because the bunching factor depends weakly on $n$ for high $n$’s (Fig. 3). Since the width of the resulting radiation is determined by the spectrum of the beam current, the long microbunch trains would generate extremely narrow spectral lines, as seen in our numerical simulation. Ultimately, one can aim at achieving the width as small as $\sim 10^{-5}$. In practice, the spectral width and phase noise of the seed laser pulse and the nonlinear energy chirp, remaining in the electron beam after the energy compression, will broaden the spectrum. Together these effects might worsen the relative width to $\sim 10^{-4}$ level. The narrow bandwidth might find applications in spectroscopy where a high spectral resolving power is required. The proposed scheme has a potential for scalability to higher beam energies, which can help mitigating detrimental effects of space charge and beam divergence. Since the radiation wavelength is defined by the harmonics of the seed wavelength, a higher beam energy may not always result in a higher photon energy. Still, it may enable a wider range of undulator parameters such as a longer period and higher strength, allowing for additional tunability of the resulting radiation.
Acknowledgments